OpenStudy (anonymous):

Write the slope-intercept form of the equation of the line that passes through (4, 2) with a slope of 3. y = 3x - 10 y = 3x + 4 y - 2 = 3 (x - 4) y = 4x + 3 wouldnt it be 2=3(4)+b?

OpenStudy (anonymous):

Yes. So b=?

OpenStudy (anonymous):

6?

OpenStudy (anonymous):

Remember point-slope form? plug in what you can and see....

OpenStudy (aripotta):

put the information in point-slope form, then change it to slope-intercept

OpenStudy (anonymous):

y-y1=m(x-x1)

OpenStudy (anonymous):

uuuuh so m=x-x1/y-y1?

OpenStudy (anonymous):

I wanted you to realize that you needed to use point slope first

OpenStudy (anonymous):

uuuuh?

OpenStudy (anonymous):

im really bad a at remember ing equation and forms what exactly is point slope?

OpenStudy (aripotta):

he already posted it

OpenStudy (aripotta):

she** sorry :/

OpenStudy (aripotta):

y-y1=m(x-x1)

OpenStudy (anonymous):

i was about to say brittni is a he?? lol

OpenStudy (aripotta):

lol yea, sorry about that brittni. i honestly didn't even look at the name until after i posted "he"

OpenStudy (aripotta):

then i realized she was a

OpenStudy (anonymous):

soooo y-4=6(x-2) i am confused

OpenStudy (aripotta):

nerp. they give a coordinate (4,2) and a slope of 3. so our y1 is gonna be 2, our x1 is gonna be 4, and our m is gonna be 3. y - 2 = 3(x - 4)

OpenStudy (anonymous):

ok thanks!:)

OpenStudy (aripotta):

wait. it's not in slope-intercept form yet. can you do that? what do you get?

OpenStudy (anonymous):

the answers are y = 3x - 10 y = 3x + 4 y - 2 = 3 (x - 4) y = 4x + 3

OpenStudy (aripotta):

yea, so which one is it?

OpenStudy (anonymous):

numero dos!

OpenStudy (aripotta):

no :/

OpenStudy (anonymous):

hay cuatro respuestas entonces es numero 2

OpenStudy (aripotta):

it's not 2

OpenStudy (anonymous):

but its the same as hwta you wrote

OpenStudy (aripotta):